# via Continuized Discrete Diffusion

### v1

@@ 3.3 Spherical Perspective Generalization
Joint training with warped images and relative embeddings reduces reliance on fixed viewpoints and improves cross-view consistency in action generation.
@@ 4 Experiment
Model LIBERO [liu2023libero] VLABench [zhang2024vlabench] ManiSkill [tao2024maniskill3] Avg.
@@
Diffusion Policy [chi2023diffusion] 78.3 92.5 68.3 50.5 - - - - - 0.0 40.0 80.0 0.0 88.0 -
MDT [reuss2024multimodal] 78.5 87.5 73.5 64.8 - - - - - - - - - - -
RDT [liu2024rdt] - - - - - - - - - 13.2 77.2 74.0 1.2 100.0 -
Dita [hou2025dita] 84.2 96.3 85.4 63.8 - - - - - - - - - - -
TraceVLA [zheng2024tracevla] 84.6 85.2 75.1 54.1 - - - - - - - - - - -
SpatialVLA [qu2025spatialvla] 88.2 89.9 78.6 55.5 - - - - - - - - - - -
OpenVLA [kim2024openvla] 84.7 88.4 79.2 53.7 - - - - - 0.0 8.0 80.0 0.0 88.0 -
Octo [mees2024octo] 78.9 85.7 84.6 51.1 - - - - - 0.0 0.0 0.0 0.0 0.0 -
$\pi_{0}$ FAST [pertsch2025fast] 96.4 96.8 88.6 60.2 46.0 42.0 26.0 30.0 20.8 0.0 80.0 52.0 0.0 92.0 52.2
$\pi_{0}$ [black2024pi_0] 96.8 98.8 95.8 85.2 54.0 48.0 16.0 6.0 7.0 4.0 60.0 48.0 0.0 100.0 51.4
$\pi_{0.5}$ [intelligence2025pi05] 95.4 98.4 97.2 89.6 24.0 18.0 36.0 20.0 6.5 8.0 56.0 56.0 4.0 92.0 50.1
@@
4.1 Training Setting
The pretrained VLM is based on Gemma [team2024gemma], which is configured as $\emph{width}=2048$, $\emph{mlp\_dim}=16384$, $\emph{depth}=18$, $\emph{num\_heads}=8$, and $\emph{head\_dim}=256$. The action expert adopts the same architecture but with a smaller configuration: $\emph{width}=1024$, $\emph{mlp\_dim}=4096$, $\emph{depth}=18$, $\emph{num\_heads}=8$, and $\emph{head\_dim}=256$. All models were fine-tuned on their corresponding datasets for 24 hours using a single NVIDIA RTX RPO6000 GPU. Each model was trained for a total of 30,00030{,}000 steps with a batch size of $32$. The learning rate followed a cosine decay schedule, implemented as CosineDecaySchedule, with a warm-up phase of 10,00010{,}000 steps, a peak learning rate of $5\times 10^{-5}$ and a final learning rate of $5\times 10^{-5}$. The optimizer is AdamW [adam2019no], configured with a gradient clipping norm of $1.0$. An exponential moving average with a decay rate of $0.999$ was applied to stabilize training. During inference, a single NVIDIA RTX 3090 GPU was used for model evaluation and deployment on the server side.
4.2 Simulation Experiment
As shown in Fig. 3, we evaluate our proposed $\mathcal{E}_{0}$ on three representative simulation benchmarks. LIBERO [liu2023libero] provides a suite of manipulation tasks designed for benchmarking robot learning. It consists of four categories—Libero-Spatial, Libero-Object, Libero-Goal, and Libero-10. Follow [liu2024rdt], we adopt a representative subset of five manipulation tasks with varying levels of difficulty—PegInsertionSide, PickCube, StackCube, PlugCharger, and PushCube—for comparative evaluation in ManiSkill [tao2024maniskill3]. VLABench [zhang2024vlabench] is specifically designed to evaluate general-purpose robotic systems built upon large multimodal models. It emphasizes language understanding, commonsense reasoning, and multi-step task planning, covering five dimensions: grid texture perception, spatial reasoning, commonsense and world knowledge, semantic understanding, and physical laws.
Analysis on LIBERO. This benchmark includes a wide range of models, many of which employ sophisticated training strategies and architectural enhancements to maximize performance. Despite the highly competitive nature of the benchmark, our model achieved the highest average performance (96%), demonstrating the effectiveness of the proposed approach even when trained with mixed data. To ensure fairness, we trained $\pi_{0}$ [black2024pi_0], $\pi_{0}$ FAST [pertsch2025fast], and $\pi_{0.5}$ [intelligence2025pi05] using the same configuration and validated them under identical settings. Experimental results show that our model exhibits more accurate grasping and is better at executing task instructions in some of the more complex scenarios. Further details are provided in the supplementary material (see Fig. 9).
[…]
@@ 4.2 Simulation Experiment
Analysis on VLABench. We observe that continuous action models [black2024pi_0] exhibit noticeable performance degradation on tasks that demand explicit reasoning or fine-grained visual discrimination. This behavior is anticipated, as [black2024pi_0] relies solely on action supervision during training, without incorporating explicit constraints for contextual reasoning. Consequently, its autoregressive backbone often struggles to maintain semantic coherence during inference. In contrast, [pertsch2025fast] shows marked improvements through its autoregressive generation framework, which enhances task-level reasoning. However, it still suffers from relatively imprecise action generation. Our proposed method, $\mathcal{E}_{0}$, achieves a balanced trade-off between these paradigms. By integrating the contextual grounding capabilities of vision–language reasoning with the fine-grained control offered by discrete diffusion, it enables robust multimodal understanding and generates precise, semantically consistent action trajectories. As shown in Fig. 4, $\pi_{0}$ [black2024pi_0] and $\pi_{0}$ FAST [pertsch2025fast] misinterpret the textual instruction and select incorrect suits, while $\pi_{0.5}$ [intelligence2025pi05] identifies the correct card but fails to grasp it precisely. In contrast, $\mathcal{E}_{0}$ both recognizes the correct suit and executes an accurate grasp, demonstrating strong semantic understanding and precise control.
@@
Figure 4: Comparison on the VLABench benchmark. In the task “pick up the spade 3”, our $\mathcal{E}_{0}$ correctly identifies and precisely grasps the target card, showing superior multimodal reasoning and control.
@@ 4.3 Real World Experiment
For real-world validation, we employ a Franka Research 3 robotic arm and evaluate our method across eight categories of manipulation tasks (see Fig. 5): pick block, press button, stack block, pull drawer, close door, pick block twice, pull drawer and put in block, and put in plate and close door. We utilize Gello [wu2024gello] to control the Franka arm and collect demonstration data. For the first five short-horizon tasks, we collect 50 trajectories per task, while for the remaining three long-horizon tasks, we collect 80 trajectories each. We train our model on a mixture of these trajectories and evaluate its performance by executing 20 trials per task. As shown in the Tab. 2, our model achieves the highest average success rate across all tasks. These results demonstrate $\mathcal{E}_{0}$’s capability to maintain semantic coherence and precise control across extended action sequences. Moreover, thanks to the strong pretrained foundation of our model, we observe impressive generalization to previously unseen configurations. For instance, in the stack block task, even when the positions of the red and green blocks are swapped—differing from the training data—the model accurately identifies their colors and completes the stacking correctly (see Fig. 5). Similarly, in the pick block twice task, while the green plate and the red and orange blocks were fixed during data collection, the model successfully handles varying placements during validation, demonstrating robust generalization and adaptability (see Fig. 14 in the supplementary material).
@@
$\pi_{0}$ [black2024pi_0] 60.0 60.0 40.0 30.0 45.0 45.0 55.0 10.0 43.1
$\pi_{0}$ FAST [pertsch2025fast] 20.0 20.0 5.0 10.0 15.0 10.0 0.0 0.0 10.0
[…]
@@ 4.4 Ablation Study
As shown in Tab. 3, both $\pi_{0}$ [black2024pi_0] and our $\mathcal{E}_{0}$ exhibit substantial performance degradation without view augmentation (average SR of 19.7% and 66.5%, respectively), indicating strong sensitivity to camera changes. Incorporating spherical view augmentation (+view) markedly improves robustness, boosting average SR to 50.8% and 83.9%. These results demonstrate that our view-aware design significantly enhances generalization to unseen or dynamic camera configurations, a key requirement for reliable real-world deployment.
@@
$\pi_{0}$ [black2024pi_0] 28.0 16.6 31.6 2.6 19.7
$\pi_{0}$ [black2024pi_0] + view 53.8 75.2 57.8 16.2 50.8 (+31.1)
[…]
@@ 4.5 Case Analysis
As shown in Fig. 7, $\mathcal{E}_{0}$ exhibits more meaningful interaction by attempting to align visual and textual cues, occasionally moving toward the correct button, though sometimes failing due to limited painting-related data and partial catastrophic forgetting during fine-tuning. Notably, $\mathcal{E}_{0}$ produces smoother, more stable motions than $\pi_{0.5}$, which often acts erratically despite achieving higher success “by chance.” We also observed inconsistent success labeling in a few cases, likely due to minor issues in the benchmark code. Overall, $\mathcal{E}_{0}$ maintains strong control quality, and future work will focus on improving its visual–language grounding to enhance task understanding.
@@ 5 Conclusion
In this work, we introduced $\mathcal{E}_{0}$, a unified VLA model that integrates discrete diffusion with AR feedback to enable fine-grained action generation. $\mathcal{E}_{0}$ formulates action prediction as a categorical diffusion process over one-hot representations. This design allows for flexible control over discretization granularity and planning horizon. Empirical results demonstrate that $\mathcal{E}_{0}$ consistently delivers superior task-level reasoning and precise action control. Collectively, our findings highlight discrete diffusion as a promising framework for advancing embodied intelligence.
[…]
Supplementary Material
6 Analysis of Discrete tokens and VLM/VLA compatibility
Lemma 1 (Discrete action tokens and hypothesis complexity).
Let $\mathcal{V}_{L}$ be a finite language token set and let the action space be represented either as a finite discrete set $\mathcal{V}_{A}$ of size $K=|\mathcal{V}_{A}|$ or as a continuous space $\mathbb{R}^{p}$. Consider hypothesis classes
$\mathcal{H}_{\mathrm{disc}}=\{g_{\theta}:X\to\Delta^{K-1}\},\hskip 28.80008pt\mathcal{H}_{\mathrm{cont}}=\{f_{\theta}:X\to\mathbb{R}^{p}\},$
where $\Delta^{K-1}$ is the probability simplex over $K$ actions. Then the following hold:
  1. 1.
    Any alignment between language tokens and discrete actions can be expressed as a mapping $\phi:\mathcal{V}_{L}\to\mathcal{V}_{A}$ of finite description length, whereas alignment to continuous actions requires a mapping $\psi:\mathcal{V}_{L}\to\mathbb{R}^{p}$ of unbounded functional complexity in general.
  2. 2.
    The hypothesis class $\mathcal{H}_{\mathrm{disc}}$ has finite VC dimension of order $\mathrm{VC}(\mathcal{H}_{\mathrm{disc}})=O(K\log K)$, yielding generalization bounds of order $O(\sqrt{K\log K/n})$ from $n$ samples. In contrast, $\mathcal{H}_{\mathrm{cont}}$ typically has infinite pseudo-dimension unless strong regularity assumptions (e.g., global Lipschitz constraints) are imposed, so its generalization bounds scale as $O\!\big(\sqrt{\mathrm{Pdim}(\mathcal{H}_{\mathrm{cont}})/n}\big)$ and may diverge in the absence of such constraints.
Consequently, discrete action tokens admit a structurally simpler and better-controlled hypothesis class than continuous actions, leading to tighter generalization guarantees and more stable semantic alignment with symbolic vision–language representations.
Proof.
Since $\mathcal{V}_{L}$ and $\mathcal{V}_{A}$ are both finite sets, any language–action alignment can be written as a function $\phi:\mathcal{V}_{L}\to\mathcal{V}_{A}$. The graph of $\phi$ is specified by a finite table of size $|\mathcal{V}_{L}|$, so $\phi$ has finite description length and can in principle be implemented exactly by a sufficiently expressive discrete model. In contrast, when actions live in $\mathbb{R}^{p}$, an alignment must be represented as a mapping $\psi:\mathcal{V}_{L}\to\mathbb{R}^{p}$. Unless one imposes additional structure (e.g., linearity or Lipschitz continuity), such a mapping may approximate arbitrarily complex functions over a continuous target space and thus has unbounded functional complexity in general. This establishes (1).
For (2), consider the discrete action case. A predictor $g_{\theta}:X\to\Delta^{K-1}$ can be viewed as a multiclass classifier over $K$ classes. Standard results in statistical learning theory state that the VC dimension of multiclass classifiers over a finite label set of size $K$ is finite and satisfies $\mathrm{VC}(\mathcal{H}_{\mathrm{disc}})=O(K\log K)$ for a broad family of architectures (e.g., linear separators, neural networks with bounded weights). This yields uniform generalization bounds of the form
$R(g_{\theta})-\hat{R}(g_{\theta})=O\!\left(\sqrt{\frac{K\log K}{n}}\right),$
where $R$ and $\hat{R}$ denote true and empirical risks, respectively, and $n$ is the number of samples.
In the continuous case, a predictor $f_{\theta}:X\to\mathbb{R}^{p}$ belongs to a real-valued function class whose complexity is captured by its pseudo-dimension $\mathrm{Pdim}(\mathcal{H}_{\mathrm{cont}})$. For general neural network or regression architectures without explicit norm or Lipschitz constraints, $\mathrm{Pdim}(\mathcal{H}_{\mathrm{cont}})$ is unbounded or grows very rapidly with the number of parameters, and can be considered effectively infinite in the worst case. Generalization bounds then take the form
$R(f_{\theta})-\hat{R}(f_{\theta})=O\!\left(\sqrt{\frac{\mathrm{Pdim}(\mathcal{H}_{\mathrm{cont}})}{n}}\right),$
which provides no meaningful guarantee when $\mathrm{Pdim}(\mathcal{H}_{\mathrm{cont}})$ is unbounded. Thus, in the absence of strong regularity assumptions, we have
$\mathrm{VC}(\mathcal{H}_{\mathrm{disc}})\ll\mathrm{Pdim}(\mathcal{H}_{\mathrm{cont}}),$
and the discrete action formulation enjoys strictly tighter sample-complexity guarantees.
Combining the finite, symbolic nature of $\mathcal{V}_{A}$ with the bounded VC dimension of $\mathcal{H}_{\mathrm{disc}}$ shows that discrete action tokens induce a hypothesis class that is both structurally aligned with token-based vision–language models and better controlled from a statistical learning perspective than continuous action heads. This proves the lemma. ∎
7 Analysis of Bayes-optimal denoiser that models correct discrete action distribution
Lemma 2 (Support Preservation of Discrete Diffusion vs. Off-Support Averaging of Continuous Diffusion).
Let the true clean action $x_{0}$ take values in a finite quantization set
$\mathcal{S}=\{b_{1},\dots,b_{K}\}\subset\mathbb{R}^{d},\hskip 28.80008pt\Pr(x_{0}=b_{k})=\pi_{k}>0.$
Consider two denoising models at timestep $t$:
  1. 1.
    A continuous diffusion denoiser $f_{\theta}(x_{t},t)$ trained with MSE, whose Bayes-optimal predictor is
    $f^{*}(x_{t},t)=\mathbb{E}[x_{0}\mid x_{t}].$
  2. 2.
    A discrete diffusion denoiser that represents actions as indices $y_{0}\in\{1,\dots,K\}$, adds Gaussian noise to the one-hot vector $e_{y_{0}}$, and predicts a categorical distribution $q_{\theta}(y\mid z_{t},t)$ optimized with cross-entropy. Its Bayes-optimal predictor is the posterior
    $q^{*}(y\mid z_{t},t)=\Pr(y_{0}=y\mid z_{t},t),$
    and its generative action is $\hat{x}_{t}=b_{\hat{y}_{t}}$ for $\hat{y}_{t}$ obtained via sampling or $\arg\max$.
Then the following statements hold:
  1. (i)
    If the posterior $\Pr(x_{0}=b_{k}\mid x_{t})$ is non-degenerate (i.e., at least two bins have positive posterior probability), then for almost all $x_{t}$,
    $f^{*}(x_{t},t)=\sum_{k=1}^{K}b_{k}\Pr(x_{0}=b_{k}\mid x_{t})\;\notin\;\mathcal{S}.$
    That is, the Bayes-optimal continuous denoiser lies off the discrete support.
  2. (ii)
    For the discrete diffusion model, the generative action always satisfies
    $\hat{x}_{t}=b_{\hat{y}_{t}}\in\mathcal{S},\hskip 28.80008pt\forall\,t.$
    Thus, discrete diffusion preserves the action support with probability 1.
Consequently, continuous diffusion exhibits mode-averaging behavior and produces off-grid actions whenever the posterior is multimodal, whereas discrete diffusion always returns a valid quantized action on $\mathcal{S}$.
Proof.
For the continuous diffusion model, the MSE-optimal denoiser is the conditional expectation
$\begin{aligned} f^{*}(x_{t},t)&=\mathbb{E}[x_{0}\mid x_{t}]=\sum_{k=1}^{K}b_{k}\,p_{k}(x_{t}),\\ p_{k}(x_{t})&=\Pr(x_{0}=b_{k}\mid x_{t})\end{aligned}.$ (9)
If the posterior is non-degenerate, there exist $i\neq j$ such that $p_{i}(x_{t}),p_{j}(x_{t})>0$ and $b_{i}\neq b_{j}$. Then $f^{*}(x_{t},t)$ is a nontrivial convex combination of distinct points in $\mathcal{S}$. Except for a measure-zero subset of $\mathbb{R}^{d}$ where exact algebraic cancellations occur, such a convex combination cannot equal any single $b_{k}$. Hence $f^{*}(x_{t},t)\notin\mathcal{S}$ for almost all $x_{t}$, establishing (i).
For the discrete diffusion model, the Bayes-optimal predictor under cross-entropy is the posterior $q^{*}(y\mid z_{t},t)$. During generation, we obtain $\hat{y}_{t}\in\{1,\dots,K\}$ either by sampling from $q^{*}$ or by taking the maximizer. Mapping the index to an action yields $\hat{x}_{t}=b_{\hat{y}_{t}}\in\mathcal{S}$. This holds deterministically for every $t$, proving (ii).
Combining the two statements shows that continuous diffusion produces off-support denoised actions whenever uncertainty is present, while discrete diffusion preserves the discrete action support by construction. ∎
8 Detailed Information for models
8.1 Architecture and Pipeline
Inference Pipeline.

### v2

@@ 3.3 Spherical Perspective Generalization
Joint training with warped images and relative embeddings reduces reliance on fixed viewpoints and improves cross-view consistency in action generation. Further details are provided in the section A.2 of the appendix.
@@ 4 Experiment
Model LIBERO (Liu et al., 2023) VLABench (Zhang et al., 2024) ManiSkill (Tao et al., 2024) Avg.
@@
Diffusion Policy (Chi et al., 2023) 78.3 92.5 68.3 50.5 - - - - - 0.0 40.0 80.0 0.0 88.0 -
MDT (Reuss et al., 2024) 78.5 87.5 73.5 64.8 - - - - - - - - - - -
RDT (Liu et al., 2024) - - - - - - - - - 13.2 77.2 74.0 1.2 100.0 -
Dita (Hou et al., 2025) 84.2 96.3 85.4 63.8 - - - - - - - - - - -
TraceVLA (Zheng et al., 2024) 84.6 85.2 75.1 54.1 - - - - - - - - - - -
SpatialVLA (Qu et al., 2025) 88.2 89.9 78.6 55.5 - - - - - - - - - - -
OpenVLA (Kim et al., 2024) 84.7 88.4 79.2 53.7 - - - - - 0.0 8.0 80.0 0.0 88.0 -
Octo (Mees et al., 2024) 78.9 85.7 84.6 51.1 - - - - - 0.0 0.0 0.0 0.0 0.0 -
$\pi_{0}$ FAST (Pertsch et al., 2025) 96.4 96.8 88.6 60.2 46.0 42.0 26.0 30.0 20.8 0.0 80.0 52.0 0.0 92.0 52.2
$\pi_{0}$ (Black et al., 2024) 96.8 98.8 95.8 85.2 54.0 48.0 16.0 6.0 7.0 4.0 60.0 48.0 0.0 100.0 51.4
$\pi_{0.5}$ (Intelligence et al., 2025) 95.4 98.4 97.2 89.6 24.0 18.0 36.0 20.0 6.5 8.0 56.0 56.0 4.0 92.0 50.1
$\mathcal{E}_{0}$ (ours) 97.2 99.4 95.0 92.2 54.0 34.0 12.0 72.0 18.8 24.0 76.0 72.0 4.0 100.0 60.8
4.1 Training Setting
The pretrained VLM is based on Gemma (Team et al., 2024), which is configured as $\emph{width}=2048$, $\emph{mlp\_dim}=16384$, $\emph{depth}=18$, $\emph{num\_heads}=8$, and $\emph{head\_dim}=256$. The action expert adopts the same architecture but with a smaller configuration: $\emph{width}=1024$, $\emph{mlp\_dim}=4096$, $\emph{depth}=18$, $\emph{num\_heads}=8$, and $\emph{head\_dim}=256$. All models were fine-tuned on their corresponding datasets for 24 hours using a single NVIDIA RTX RPO6000 GPU. Each model was trained for a total of 30,00030{,}000 steps with a batch size of $32$. The learning rate followed a cosine decay schedule, implemented as CosineDecaySchedule, with a warm-up phase of 1,0001{,}000 steps, a peak learning rate of $5\times 10^{-5}$ and a final learning rate of $5\times 10^{-6}$. The optimizer is AdamW (Adam et al., 2019), configured with a gradient clipping norm of $1.0$. An exponential moving average with a decay rate of $0.999$ was applied to stabilize training. During inference, a single NVIDIA RTX 3090 GPU was used for model evaluation and deployment on the server side.
While our model adopts the same backbone network as $\pi_{0}$ FAST (Pertsch et al., 2025), $\pi_{0}$ (Black et al., 2024), and $\pi_{0.5}$ (Intelligence et al., 2025), the action modeling in these methods relies on autoregressive or diffusion-based action heads. Consequently, we consider them as baselines and retrain them using the same experimental setup to ensure a fair comparison and to control for variations in training settings and environments.
4.2 Simulation Experiment
As shown in Fig. 3, we evaluate our proposed $\mathcal{E}_{0}$ on three representative simulation benchmarks. LIBERO (Liu et al., 2023) provides a suite of manipulation tasks designed for benchmarking robot learning. It consists of four categories—Libero-Spatial, Libero-Object, Libero-Goal, and Libero-10. Following RDT (Liu et al., 2024), we adopt a representative subset of five manipulation tasks with varying levels of difficulty—PegInsertionSide, PickCube, StackCube, PlugCharger, and PushCube—for comparative evaluation in ManiSkill (Tao et al., 2024). VLABench (Zhang et al., 2024) is specifically designed to evaluate general-purpose robotic systems built upon large multimodal models. It emphasizes language understanding, commonsense reasoning, and multi-step task planning, covering five dimensions: grid texture perception, spatial reasoning, commonsense and world knowledge, semantic understanding, and physical laws.
Analysis on LIBERO. This benchmark includes a wide range of models, many of which employ sophisticated training strategies and architectural enhancements to maximize performance. Despite the highly competitive nature of the benchmark, our model achieved the highest average performance (96%), demonstrating the effectiveness of the proposed approach even when trained with mixed data. To ensure fairness, we trained $\pi_{0}$, $\pi_{0}$ FAST, and $\pi_{0.5}$ using the same configuration and validated them under identical settings. Experimental results show that our model exhibits more accurate grasping and is better at executing task instructions in some of the more complex scenarios. Further details are provided in the section D of the appendix.
@@ 4.2 Simulation Experiment
Analysis on VLABench. We observe that continuous action models exhibit noticeable performance degradation on tasks that demand explicit reasoning or fine-grained visual discrimination. This behavior is anticipated, as $\pi_{0}$ relies solely on action supervision during training, without incorporating explicit constraints for contextual reasoning. Consequently, its autoregressive backbone often struggles to maintain semantic coherence during inference. In contrast, $\pi_{0}$ FAST shows marked improvements through its autoregressive generation framework, which enhances task-level reasoning. However, it still suffers from relatively imprecise action generation. Our proposed method, $\mathcal{E}_{0}$, achieves a balanced trade-off between these paradigms. By integrating the contextual grounding capabilities of vision–language reasoning with the fine-grained control offered by discrete diffusion, it enables robust multimodal understanding and generates precise, semantically consistent action trajectories. As shown in Fig. 4, $\pi_{0}$ and $\pi_{0}$ FAST misinterpret the textual instruction and select incorrect suits, while $\pi_{0.5}$ identifies the correct card but fails to grasp it precisely. In contrast, $\mathcal{E}_{0}$ both recognizes the correct suit and executes an accurate grasp, demonstrating strong semantic understanding and precise control.
@@
Figure 4: Comparison on the VLABench benchmark. In the task “pick up the spade 3”, our $\mathcal{E}_{0}$ correctly identifies and precisely grasps the target card, showing superior multimodal reasoning and control ability.
@@ 4.3 Real World Experiment
For real-world validation, we employ a Franka Research 3 robotic arm and evaluate our method across eight categories of manipulation tasks (see Fig. 5). We utilize Gello (Wu et al., 2024) to control the Franka arm and collect demonstration data. For the short-horizon tasks, we collect 50 trajectories per task, while for the long-horizon tasks, we collect 80 trajectories each. We train our model on a mixture of these trajectories and evaluate its performance by executing 20 trials per task. As shown in the Tab. 2, our model achieves the highest average success rate across all tasks. These results demonstrate $\mathcal{E}_{0}$’s capability to maintain semantic coherence and precise control across extended action sequences. Moreover, we observe impressive generalization to previously unseen configurations. For instance, in the stack block task, even when the positions of the red and green blocks are swapped—differing from the training data—the model accurately identifies their colors and completes the stacking correctly (see Fig. 5). Similarly, in the pick block twice task, while the green plate and the red and orange blocks were fixed during data collection, the model successfully handles varying placements during validation, demonstrating robust generalization and adaptability. Further details are provided in the section H of the appendix.
@@
$\pi_{0}$ (Black et al., 2024) 60.0 60.0 40.0 30.0 45.0 45.0 55.0 10.0 43.1
$\pi_{0}$ FAST (Pertsch et al., 2025) 20.0 20.0 5.0 10.0 15.0 10.0 0.0 0.0 10.0
$\mathcal{E}_{0}$ (ours) 75.0 70.0 50.0 35.0 35.0 60.0 40.0 20.0 45.6
@@ 4.4 Ablation Study
As shown in Tab. 3, both $\pi_{0}$ and our $\mathcal{E}_{0}$ exhibit substantial performance degradation without view augmentation (average SR of 19.7% and 66.5%, respectively), indicating strong sensitivity to camera changes. Incorporating spherical view augmentation (+view) markedly improves robustness, boosting average SR to 50.8% and 83.9%. These results demonstrate that our view-aware design significantly enhances generalization to unseen or dynamic camera configurations, a key requirement for reliable real-world deployment.
@@
$\pi_{0}$ (Black et al., 2024) 28.0 16.6 31.6 2.6 19.7
$\pi_{0}$ (Black et al., 2024) + view 53.8 75.2 57.8 16.2 50.8 (+31.1)
$\mathcal{E}_{0}$ (Ours) 74.4 85.2 77.4 28.8 66.5
$\mathcal{E}_{0}$ (Ours) + view 92.0 96.8 83.0 63.8 83.9 (+22.6)
@@ 4.5 Case Analysis
As shown in Fig. 7, $\mathcal{E}_{0}$ exhibits more meaningful interaction by attempting to align visual and textual cues, occasionally moving toward the correct button, though sometimes failing due to limited painting-related data and partial catastrophic forgetting during fine-tuning. Notably, $\mathcal{E}_{0}$ produces smoother, more stable motions than $\pi_{0.5}$, which often acts erratically despite achieving higher success “by chance.” We also observed inconsistent success labeling in a few cases, likely due to minor issues in the benchmark code. Overall, $\mathcal{E}_{0}$ maintains strong control quality, and future work will focus on improving its visual–language grounding to enhance task understanding. Further details are provided in the section F of the appendix.
@@ 5 Conclusion
In this work, we revisited action modeling in Vision–Language–Action systems from a structural perspective. We showed that the multi-peaked nature of action distributions, the token-based reasoning of pretrained VLM/VLA backbones, and the effective finite resolution of real-world control jointly motivate discrete action modeling. Based on these insights, we proposed $\mathcal{E}_{0}$, a Tweedie discrete diffusion framework for principled and fine-grained action generation. Our results demonstrate that discrete diffusion provides a robust and generalizable foundation for action modeling in VLA systems.
Impact Statement
This work aims to advance Vision–Language–Action models for robotic manipulation by improving the robustness and generalization of action generation through a discrete diffusion-based method. As a methodological contribution, our approach is intended to support research on learning-based robotic control rather than direct deployment in safety-critical or autonomous decision-making systems.
Potential societal impacts primarily arise in real-world robotic applications, where improperly generated actions could lead to unsafe behavior or unintended interactions with the physical environment. These risks are not unique to our method and are common to learning-based robotic control in general. In practice, such risks can be mitigated through standard safeguards, including deployment in controlled environments, human supervision, simulation-based validation, and existing physical safety mechanisms on robotic platforms. We do not anticipate new ethical concerns beyond those already well established in the field of robot learning.
Appendix A Implementation and Model Details
A.1 Architecture and Pipeline